OpenStudy (anonymous):

Find the solutions of the equation. 1/2x^2 - x + 5 = 0

OpenStudy (anonymous):

can someon eexplain to me who to do this

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\] where you already have an equation of the form ax^ + bx + c = 0. a=1/2, b=-1, and c=5. Form equation from substitutions.

OpenStudy (anonymous):

Hint: Concentrate first on the expression b^2 - 4ac, which is the determinant. If this expression is less than zero, the equation has no real number solutions. If the equation has no real number solutions, it still will have complex number solutions given from the equation for "x" in my first post. You may or may not have gotten to the point in your studies where your class is dealing with complex numbers. Try to determine if your teacher is looking for you to supply complex number solutions or only real number solutions.